OpenStudy (anonymous):

what percent of 64 is 20?

OpenStudy (anonymous):

X*64=20

OpenStudy (anonymous):

Solve for X

OpenStudy (anonymous):

It will be a decimal number so will have to convert it from decimal to percent.

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

First solve for x and tell me what you got.

OpenStudy (anonymous):

3.2

OpenStudy (anonymous):

Nope that is wrong. X*64=20 X=20/64 ----divide 64 to both sides What is 20/64

OpenStudy (anonymous):

.3125

OpenStudy (anonymous):

Great!!! But you want percentage All you do is move the decimal point two times tot the right. Then just stick a percentage sign in the end.

OpenStudy (anonymous):

.3125 = 31.25%

OpenStudy (anonymous):

ooooh okay,,, thanks